# The barrier Ramsey theorem

### v2

@@ 7. The Upper Bound
Starting from $\mathcal{A}$ and $\mathcal{C}$ we define sets of $\mathsf{SD}$-barriers $(\mathcal{F}_{s})_{s\in T(\mathcal{A})}$ and $(\mathcal{E}_{s})_{s\in T(\mathcal{A})}$ where $\operatorname{base}(\mathcal{E}_{s})=\operatorname{base}(\mathcal{F}_{s})=\operatorname{base}(\mathcal{A})\setminus\{0,\ldots,\max s\}$. If $s\in\mathcal{A}$ then we define $\mathcal{F}_{s}=\mathcal{C}$ restricted to $\operatorname{base}(\mathcal{A})\setminus\{0,\ldots,\max s\}$. If $s\in T(\mathcal{A})\setminus\mathcal{A}$, assume that for each $n>\max s$ we have defined $\mathcal{F}_{s{}^{\smallfrown}\langle n\rangle}$. Let $\mathcal{E}_{s{}^{\smallfrown}\langle n\rangle}$ be a $\mathsf{SD}$-barrier (whose existence is ensured by the assumption on Theorem 7.1 and by the fact that $1+0pt(\mathcal{C}_{n})<\omega^{\delta}$ for each $n$) with $\operatorname{base}(\mathcal{E}_{s{}^{\smallfrown}\langle n\rangle})=\operatorname{base}(\mathcal{F}_{s{}^{\smallfrown}\langle n\rangle})$, $0pt(\mathcal{E}_{s{}^{\smallfrown}\langle n\rangle})\leq\varphi_{\log 0pt(\mathcal{C}_{n})}(0pt(\mathcal{F}_{s{}^{\smallfrown}\langle n\rangle})\mathbin{{\times}\!\!\!\!{\times}}k^{(2^{|s|})})$ and $\mathcal{E}_{s{}^{\smallfrown}\langle n\rangle}\rightarrow(\mathcal{F}_{s{}^{\smallfrown}\langle n\rangle})^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}_{n}}_{k^{(2^{|s|})}}$. We define $\mathcal{F}_{s}$ as the set $\{t\subseteq\operatorname{base}(\mathcal{A}):t^{-}\in\mathcal{E}_{s{}^{\smallfrown}\langle\min t\rangle}\}$. Then $\mathcal{F}_{s}$ is clearly a block with $\operatorname{base}(\mathcal{F}_{s})=\operatorname{base}(\mathcal{A})\setminus\{0,\ldots,\max s\}$. By Lemma 2.19 and Corollary 2.32 we may assume that $\mathcal{F}_{s}$ is a $\mathsf{SD}$-barrier with same height and same base.
We claim that for each $s\in T(\mathcal{A})$, $0pt(\mathcal{F}_{s})\leq\varphi_{\delta}(0pt(\mathcal{A}_{s}))$. We proceed by induction on $0pt(\mathcal{A}_{s})$. If $s\in\mathcal{A}$ then $\mathcal{F}_{s}=\mathcal{C}$ and $0pt(\mathcal{F}_{s})=\omega^{\delta}\leq\varphi_{\delta}(0)=\varphi_{\delta}(0pt(\mathcal{A}_{s}))$. Now suppose that $s\in T(\mathcal{A})\setminus\mathcal{A}$ and the claim holds for all $s{}^{\smallfrown}\langle n\rangle\in T(\mathcal{A})$. By construction, $0pt(\mathcal{E}_{s{}^{\smallfrown}\langle n\rangle})\leq\varphi_{\log 0pt(\mathcal{C}_{n})}(0pt(\mathcal{F}_{s{}^{\smallfrown}\langle n\rangle})\mathbin{{\times}\!\!\!\!{\times}}k^{(2^{|s|})})$. It follows that
@@
Suppose we are at stage $n$ and that we have already defined $f_{n-1}$ and $F_{n-1}$ satisfying clauses $1\div 5$. Let $f_{n}=\min F_{n-1}$ (so that clause $2$ is satisfied) and let $E_{n}=F_{n-1}^{-}$. Let $c^{\prime}_{n}\colon[E_{n}]^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}_{f_{n}}}\rightarrow k^{(2^{n})}$ be the function that maps $e\in[E_{n}]^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}_{f_{n}}}$ to (a code for) the function $u_{n-1}]^{\mathcal{D}}\rightarrow k$ defined by $d\mapsto c(d{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}e)$. Since $u_{n-1}$ has $2^{n}$ subsets, then $u_{n-1}]^{\mathcal{D}}$ has at most $2^{n}$ elements. Therefore there are at most $k^{(2^{n})}$ functions $u_{n-1}]^{\mathcal{D}}\rightarrow k$ and so $c^{\prime}_{n}$ is well defined. We know that $F_{n-1}$ is $(1\oplus\mathcal{F}_{u_{n-1}})$-large and so $F_{n-1}^{*}$ is $\mathcal{F}_{u_{n-1}}$-large. It follows that $E_{n}^{*}=F_{n-1}^{*-}$ is $\mathcal{E}_{u_{n}}$-large and hence $E_{n}$ is $(\operatorname{\mymathbb{1}}\oplus\mathcal{E}_{u_{n}})$-large. By construction $\mathcal{E}_{u_{n}}\rightarrow(\mathcal{F}_{u_{n}})^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}_{f_{n}}}_{k^{(2^{n})}}$ and so there is a $(\operatorname{\mymathbb{1}}\oplus\mathcal{C}_{f_{n}}\oplus\mathcal{F}_{u_{n}})$-large subset of $E_{n}$ which is $c^{\prime}_{n}$-homogeneous. We call such set $F_{n}$ and notice that $F_{n}$ is $(\operatorname{\mymathbb{1}}\oplus\mathcal{F}_{u_{n}})$-large (hence clauses $3$ and $4$ are satisfied). Since $F_{n}\subseteq E_{n}$, $f_{n}<F_{n}$ and clause $1$ holds too. We are left to prove that clause $5$ is satisfied. Let $t_{1},t_{2}\in[s]^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}\oplus\mathcal{D}}$ and let $t_{1}^{\prime}$ and $t_{2}^{\prime}$ be their $(\operatorname{\mymathbb{1}}\oplus\mathcal{D})$-size prefixes. Moreover assume that $t_{1}^{\prime}=t_{2}^{\prime}\subseteq u_{n}$ and $t_{1}\setminus t_{1}^{\prime},t_{2}\setminus t_{2}^{\prime}\subseteq F_{n}$. If $t_{1}^{\prime}=t_{2}^{\prime}\subseteq u_{n-1}$ then since $F_{n}\subseteq F_{n-1}$, by clause $5$ of stage n1n-1 we are done. If not then $t_{1}^{\prime*}=t_{2}^{\prime*}\in[u_{n-1}]^{\mathcal{D}}$ and $\max t_{1}^{\prime}=\max t_{n}^{\prime}=f_{n}$. Since $F_{n}$ is $c^{\prime}_{n}$-homogeneous, then $d\mapsto c(d{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}t_{1}\setminus t_{1}^{\prime})$ and $d\mapsto c(d{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}t_{2}\setminus t_{2}^{\prime})$ are the same function (notice that $t_{1}\setminus t_{1}^{\prime}$ and $t_{2}\setminus t_{2}^{\prime}$ are $(\operatorname{\mymathbb{1}}\oplus\mathcal{C}_{f_{n}})$-size). Therefore $c(t_{1})=c(t_{1}^{\prime*}{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}t_{1}\setminus t_{1}^{\prime})=c(t_{2}^{\prime*}{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}t_{2}\setminus t_{2}^{\prime})=c(t_{2})$.
@@
The construction stops at a stage $m$ such that $u_{m}\in\mathcal{A}$. At this stage $\mathcal{F}_{u_{m}}=\mathcal{C}$ by definition and so $F_{m}$ is $(\operatorname{\mymathbb{1}}\oplus\mathcal{C})$-large. Let $w\sqsubseteq F_{m}$ be $(\operatorname{\mymathbb{1}}\oplus\mathcal{C})$-size and let $f_{m+1}=\min w$. The set $v=u_{m}{}^{\smallfrown}w$ is $(\operatorname{\mymathbb{1}}\oplus\mathcal{C}\oplus\mathcal{A})$-size. Let $t_{1},t_{2}\in[v]^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}\oplus\mathcal{D}}$ and let $t_{1}^{\prime}$ and $t_{2}^{\prime}$ be their $(\operatorname{\mymathbb{1}}\oplus\mathcal{D})$-size prefixes. Moreover assume that $t_{1}^{\prime}=t_{2}^{\prime}$. Suppose first that $\max t_{1}^{\prime}=\max t_{2}^{\prime}\geq f_{m+1}$ then it must be $t_{1}\setminus t_{1}^{\prime*},t_{2}\setminus t_{2}^{\prime*}\in[w]^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}}$ and since $\operatorname{\mymathbb{1}}\oplus\mathcal{C}$ is a barrier we get $t_{1}\setminus t_{1}^{\prime*}=w=t_{2}\setminus t_{2}^{\prime*}$. Hence $t_{1}=t_{2}$ and $c(t_{1})=c(t_{2})$. Suppose now that $\max t_{1}^{\prime}=\max t_{2}^{\prime}=f_{n}$ for some $n\leq m$ and so $t_{1}^{\prime*}=t_{2}^{\prime*}\in[u_{n-1}]^{\mathcal{D}}$. Since $F_{n}$ is $c^{\prime}_{n}$-homogeneous and $\langle f_{n+1},\ldots,f_{m}\rangle{}^{\smallfrown}w\subseteq F_{n}$, then $d\mapsto c(d{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}t_{1}\setminus t_{1}^{\prime})$ and $d\mapsto c(d{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}t_{2}\setminus t_{2}^{\prime})$ are the same function. Therefore
@@
We proceed by induction on $\gamma$. If $\gamma=0$ then $\mathcal{B}=\mathcal{A}$ works. If $\gamma>0$ the induction hypothesis and the modularity of the proof that Theorem 7.4 implies Theorem 7.1 yield that Theorem 7.4 holds for every $\gamma^{\prime}<\gamma$.
The case $\gamma=\omega^{\delta}$ is proved in Lemmas 7.5 and 7.6.
@@ 8. Nestedness of the System
  • Suppose $\delta=\varphi_{\delta_{0}}(\alpha)$ with $\delta>\alpha>0$ and $\delta_{0}>0$. We consider three subcases depending on the value of $\alpha$:
    • when $\alpha=0$ we must have $\delta_{0}>\Gamma_{0}$; by inductive hypothesis $\delta_{0}[1]\geq\Gamma_{0}$ and so $\delta[1]=\varphi^{2}_{\delta_{0}[1]}(0)>\Gamma_{0}$;
    • when $0<\alpha<\Gamma_{0}$ we must have $\delta_{0}\geq\Gamma_{0}$; then we obtain $\delta[1]=\varphi^{2}_{\delta_{0}[1]}(\varphi_{\delta_{0}}(\alpha[1])+1)>\Gamma_{0}$;
    • finally, if $\alpha>\Gamma_{0}$, by inductive hypothesis we have $\alpha[1]\geq\Gamma_{0}$, and then $\delta[1]=\varphi^{2}_{\delta_{0}[1]}(\varphi_{\delta}(\alpha[1])+1)>\Gamma_{0}$.
  • If $\delta=\Gamma_{\xi}$ then $\delta[1]=\varphi_{\varphi_{\Gamma_{\xi\lceil n\rceil}+1}(0)}(0)>\Gamma_{0}$.∎
@@
  • If $\beta=\varphi_{\xi}(0)$ then $\xi>\gamma$ and by inductive hypothesis $\xi[1]\geq\gamma$. Hence $\beta[1]=\varphi^{2}_{\xi[1]}(0)>\gamma$.
  • If $\beta=\varphi_{\xi}(\alpha)$ with $0<\alpha\leq\gamma$ then it must be $\xi\geq\gamma$. Hence $\beta[1]=\varphi^{2}_{\xi[1]}(\varphi_{\xi}(\alpha[1])+1)\geq\varphi^{2}_{\xi[1]}(\varphi_{\gamma}(\alpha[1])+1)>\gamma$.
  • If $\beta=\varphi_{\xi}(\alpha)$ with $\alpha>\gamma$ then by inductive hypothesis $\alpha[1]\geq\gamma$. Hence $\beta[1]=\varphi^{2}_{\xi[1]}(\varphi_{\xi}(\alpha[1])+1)>\varphi^{2}_{\xi[1]}(\gamma)\geq\gamma$. ∎
@@
Let $\gamma$ and $\delta$ be ordinals such that $\gamma$ is not of the form $\varphi_{\gamma_{0}}(0)$ for any $\gamma_{0}$. If $\gamma>\varphi_{\delta}(0)$ then $\gamma[1]\geq\varphi_{\delta}(0)$.
@@
If $\gamma$ is decomposable or $\delta=0$ the thesis is trivial. We can assume $\gamma=\varphi_{\xi}(\gamma_{0})>\gamma_{0}>0$. Case $\delta=1$ was proved in Lemma 8.2. Case $\xi<\delta$ follows by Lemma 8.3. Finally, if $\xi\geq\delta$ then $\gamma[1]=\varphi^{2}_{\xi[1]}(\varphi_{\xi}(\gamma_{0}[1])+1)\geq\varphi^{2}_{\xi[1]}(\varphi_{\delta}(\gamma_{0}[1])+1)>\varphi_{\delta}(0)$. ∎
@@
recall that $\gamma[n]=\omega^{\gamma_{0}}+\gamma_{1}[n]$. To prove 1. assume towards a contradiction that $\gamma[n]<\beta[n]<\gamma<\beta$ and consider different subcases depending on the form of $\beta$.
@@
    $\beta=\Gamma_{\xi}$. Let $\delta_{n+1}=\Gamma_{\xi\lceil n\rceil}+1$ and $\delta_{i}=\varphi_{\delta_{i+1}}(0)$ for $i\leq n$. Thus $\Gamma_{\xi}[n]=\varphi_{\delta_{0}}(0)$. Then our assumption $\gamma[n]<\beta[n]<\gamma<\beta$ yields $\gamma_{0}[n]<\delta_{0}<\gamma_{0}<\beta$. We claim that $\gamma_{0}=\varphi_{\gamma_{1}}(0)$ for some $\gamma_{1}$. In fact, if $\gamma_{0}$ were decomposable then, since $\delta_{0}$ is indecomposable, $\gamma_{0}[n]\geq\delta_{0}$, a contradiction. If instead $\gamma_{0}=\varphi_{\gamma_{1}}(\gamma_{2})>\gamma_{2}>0$ then, since $\delta_{0}=\varphi_{\delta_{1}}(0)$, Lemma 8.4 implies $\gamma_{0}[n]\geq\delta_{0}$. This proves the claim. Notice that it must be $\gamma_{0}>\gamma_{1}>0$. The claim and the chain of inequalities $\gamma_{0}[n]<\delta_{0}<\gamma_{0}<\beta$ yields $\gamma_{1}[n]<\delta_{1}<\gamma_{1}<\beta$. Iterating $n+1$ times we obtain for each $i\leq n+1$ an ordinal $\gamma_{i}$ such that $\gamma_{i-1}=\varphi_{\gamma_{i}}(0)$ and $\gamma_{i}[n]<\delta_{i}<\gamma_{i}<\beta$. In particular $\gamma_{n+1}[n]<\delta_{n+1}<\gamma_{n+1}<\beta$. Since $\gamma_{n+1}$ is not a $\Gamma$ ordinal and $\gamma_{n+1}>\delta_{0}=\Gamma_{\xi\lceil n\rceil}+1$, Lemma 8.3 implies $\gamma_{n+1}[n]\geq\Gamma_{\xi\lceil n\rceil}$. Since $\Gamma_{\xi\lceil n\rceil}+1>\gamma_{n+1}[n]$ it must be $\gamma_{n+1}[n]=\Gamma_{\xi\lceil n\rceil}$. Lemma 8.5 implies that $\gamma_{n+1}=\Gamma_{\xi\lceil n\rceil}+1$, contradicting $\Gamma_{\xi\lceil n\rceil}+1=\delta_{n+1}<\gamma_{n+1}$.

### v3

@@ 7. The Upper Bound
Starting from $\mathcal{A}$ and $\mathcal{C}$ we define sets of $\mathsf{SD}$-barriers $(\mathcal{F}_{s})_{s\in T(\mathcal{A})}$ and $(\mathcal{E}_{s})_{s\in T(\mathcal{A})}$ where $\operatorname{base}(\mathcal{E}_{s})=\operatorname{base}(\mathcal{F}_{s})=\operatorname{base}(\mathcal{A})\setminus\{0,\ldots,\max s\}$. If $s\in\mathcal{A}$ then we define $\mathcal{F}_{s}=\mathcal{C}$ restricted to $\operatorname{base}(\mathcal{A})\setminus\{0,\ldots,\max s\}$ (the restriction is still a $\mathsf{SD}$-barrier by Remark 2.33). If $s\in T(\mathcal{A})\setminus\mathcal{A}$, assume that for each $n>\max s$ we have defined $\mathcal{F}_{s{}^{\smallfrown}\langle n\rangle}$. Let $\mathcal{E}_{s{}^{\smallfrown}\langle n\rangle}$ be a $\mathsf{SD}$-barrier (whose existence is ensured by the assumption on Theorem 7.1 and by the fact that $1+0pt(\mathcal{C}_{n})<\omega^{\delta}$ for each $n$) with $\operatorname{base}(\mathcal{E}_{s{}^{\smallfrown}\langle n\rangle})=\operatorname{base}(\mathcal{F}_{s{}^{\smallfrown}\langle n\rangle})$, $0pt(\mathcal{E}_{s{}^{\smallfrown}\langle n\rangle})\leq\varphi_{\log 0pt(\mathcal{C}_{n})}(0pt(\mathcal{F}_{s{}^{\smallfrown}\langle n\rangle})\mathbin{{\times}\!\!\!\!{\times}}k^{(2^{|s|})})$ and $\mathcal{E}_{s{}^{\smallfrown}\langle n\rangle}\rightarrow(\mathcal{F}_{s{}^{\smallfrown}\langle n\rangle})^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}_{n}}_{k^{(2^{|s|})}}$. We define $\mathcal{F}_{s}$ as the set $\{t\subseteq\operatorname{base}(\mathcal{A}):t^{-}\in\mathcal{E}_{s{}^{\smallfrown}\langle\min t\rangle}\}$. Then $\mathcal{F}_{s}$ is clearly a block with $\operatorname{base}(\mathcal{F}_{s})=\operatorname{base}(\mathcal{A})\setminus\{0,\ldots,\max s\}$. By Lemma 2.19 and Corollary 2.32 we may assume that $\mathcal{F}_{s}$ is a $\mathsf{SD}$-barrier with same height and same base.
We claim that for each $s\in T(\mathcal{A})$, $0pt(\mathcal{F}_{s})\leq\varphi_{\delta}(0pt(\mathcal{A}_{s}))$. We proceed by induction on $0pt(\mathcal{A}_{s})$. If $s\in\mathcal{A}$ then $\mathcal{F}_{s}$ is the restriction of $\mathcal{C}$ to a final segment of its base and by Lemma 2.23 $0pt(\mathcal{F}_{s})=\omega^{\delta}\leq\varphi_{\delta}(0)=\varphi_{\delta}(0pt(\mathcal{A}_{s}))$. Now suppose that $s\in T(\mathcal{A})\setminus\mathcal{A}$ and the claim holds for all $s{}^{\smallfrown}\langle n\rangle\in T(\mathcal{A})$. By construction, $0pt(\mathcal{E}_{s{}^{\smallfrown}\langle n\rangle})\leq\varphi_{\log 0pt(\mathcal{C}_{n})}(0pt(\mathcal{F}_{s{}^{\smallfrown}\langle n\rangle})\mathbin{{\times}\!\!\!\!{\times}}k^{(2^{|s|})})$. It follows that
@@
Suppose we are at stage $n$ and that we have already defined $f_{n-1}$ and $F_{n-1}$ satisfying clauses $1\div 5$. Let $f_{n}=\min F_{n-1}$ (so that clause $2$ is satisfied) and let $E_{n}=F_{n-1}^{-}$. Let $c^{\prime}_{n}\colon[E_{n}]^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}_{f_{n}}}\rightarrow k^{(2^{n})}$ be the function that maps $e\in[E_{n}]^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}_{f_{n}}}$ to (a code for) the function $u_{n-1}]^{\mathcal{D}}\rightarrow k$ defined by $d\mapsto c(d{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}e)$. Since $u_{n-1}$ has $2^{n}$ subsets, then $u_{n-1}]^{\mathcal{D}}$ has at most $2^{n}$ elements. Therefore there are at most $k^{(2^{n})}$ functions $u_{n-1}]^{\mathcal{D}}\rightarrow k$ and so $c^{\prime}_{n}$ is well defined. We know that $F_{n-1}$ is $(1\oplus\mathcal{F}_{u_{n-1}})$-large and so $F_{n-1}^{*}$ is $\mathcal{F}_{u_{n-1}}$-large. It follows that $E_{n}^{*}=F_{n-1}^{*-}$ is $\mathcal{E}_{u_{n}}$-large and hence $E_{n}$ is $(\operatorname{\mymathbb{1}}\oplus\mathcal{E}_{u_{n}})$-large. By construction $\mathcal{E}_{u_{n}}\rightarrow(\mathcal{F}_{u_{n}})^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}_{f_{n}}}_{k^{(2^{n})}}$ and so there is a $(\operatorname{\mymathbb{1}}\oplus\mathcal{C}_{f_{n}}\oplus\mathcal{F}_{u_{n}})$-large subset of $E_{n}$ which is $c^{\prime}_{n}$-homogeneous. We call such set $F_{n}$ and notice that $F_{n}$ is $(\operatorname{\mymathbb{1}}\oplus\mathcal{F}_{u_{n}})$-large (hence clauses $3$ and $4$ are satisfied). Since $F_{n}\subseteq E_{n}$, $f_{n}<F_{n}$ and clause $1$ holds too. We are left to prove that clause $5$ is satisfied. Let $t_{1},t_{2}\in[s]^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}\oplus\mathcal{D}}$ and let $t_{1}^{\prime}$ and $t_{2}^{\prime}$ be their $(\operatorname{\mymathbb{1}}\oplus\mathcal{D})$-size prefixes. Moreover assume that $t_{1}^{\prime}=t_{2}^{\prime}\subseteq u_{n}$ and $t_{1}\setminus t_{1}^{\prime},t_{2}\setminus t_{2}^{\prime}\subseteq F_{n}$. If $t_{1}^{\prime}=t_{2}^{\prime}\subseteq u_{n-1}$ then since $F_{n}\subseteq F_{n-1}$, by clause $5$ of stage n1n-1 we are done. If not then $t_{1}^{\prime*}=t_{2}^{\prime*}\in[u_{n-1}]^{\mathcal{D}}$ and $\max t_{1}^{\prime}=\max t_{2}^{\prime}=f_{n}$. Since $F_{n}$ is $c^{\prime}_{n}$-homogeneous, then $d\mapsto c(d{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}t_{1}\setminus t_{1}^{\prime})$ and $d\mapsto c(d{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}t_{2}\setminus t_{2}^{\prime})$ are the same function (notice that $t_{1}\setminus t_{1}^{\prime}$ and $t_{2}\setminus t_{2}^{\prime}$ are $(\operatorname{\mymathbb{1}}\oplus\mathcal{C}_{f_{n}})$-size). Therefore $c(t_{1})=c(t_{1}^{\prime*}{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}t_{1}\setminus t_{1}^{\prime})=c(t_{2}^{\prime*}{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}t_{2}\setminus t_{2}^{\prime})=c(t_{2})$.
@@
The construction stops at a stage $m$ such that $u_{m}\in\mathcal{A}$. At this stage $\mathcal{F}_{u_{m}}$ is the restriction of $\mathcal{C}$ to a final segment of its base by definition and so $F_{m}$ is $(\operatorname{\mymathbb{1}}\oplus\mathcal{C})$-large. Let $w\sqsubseteq F_{m}$ be $(\operatorname{\mymathbb{1}}\oplus\mathcal{C})$-size and let $f_{m+1}=\min w$. The set $v=u_{m}{}^{\smallfrown}w$ is $(\operatorname{\mymathbb{1}}\oplus\mathcal{C}\oplus\mathcal{A})$-size. Let $t_{1},t_{2}\in[v]^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}\oplus\mathcal{D}}$ and let $t_{1}^{\prime}$ and $t_{2}^{\prime}$ be their $(\operatorname{\mymathbb{1}}\oplus\mathcal{D})$-size prefixes. Moreover assume that $t_{1}^{\prime}=t_{2}^{\prime}$. Suppose first that $\max t_{1}^{\prime}=\max t_{2}^{\prime}\geq f_{m+1}$ then it must be $t_{1}\setminus t_{1}^{\prime*},t_{2}\setminus t_{2}^{\prime*}\in[w]^{\operatorname{\mymathbb{1}}\oplus\mathcal{C}}$ and since $\operatorname{\mymathbb{1}}\oplus\mathcal{C}$ is a barrier we get $t_{1}\setminus t_{1}^{\prime*}=w=t_{2}\setminus t_{2}^{\prime*}$. Hence $t_{1}=t_{2}$ and $c(t_{1})=c(t_{2})$. Suppose now that $\max t_{1}^{\prime}=\max t_{2}^{\prime}=f_{n}$ for some $n\leq m$ and so $t_{1}^{\prime*}=t_{2}^{\prime*}\in[u_{n-1}]^{\mathcal{D}}$. Since $F_{n}$ is $c^{\prime}_{n}$-homogeneous and $\langle f_{n+1},\ldots,f_{m}\rangle{}^{\smallfrown}w\subseteq F_{n}$, then $d\mapsto c(d{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}t_{1}\setminus t_{1}^{\prime})$ and $d\mapsto c(d{}^{\smallfrown}\langle f_{n}\rangle{}^{\smallfrown}t_{2}\setminus t_{2}^{\prime})$ are the same function. Therefore
@@
We proceed by induction on $\gamma$. If $\gamma=0$ then $\mathcal{B}=\mathcal{A}$ works.
If $\gamma>0$ the induction hypothesis and the modularity of the proof that Theorem 7.4 implies Theorem 7.1 yield that Theorem 7.1 holds for every $\gamma^{\prime}<\gamma$. Then case $\gamma=\omega^{\delta}$ is proved in Lemmas 7.5 and 7.6.
@@ 8. Nestedness of the System
  • Suppose $\delta=\varphi_{\delta_{0}}(\alpha)$ with $\delta>\alpha>0$ and $\delta_{0}>0$. We consider subcases depending on the value of $\alpha$:
    • when $0<\alpha\leq\Gamma_{0}$ we must have $\delta_{0}\geq\Gamma_{0}$ as otherwise $\delta\leq\Gamma_{0}$, a contradiction; then we obtain $\delta[1]=\varphi^{2}_{\delta_{0}[1]}(\varphi_{\delta_{0}}(\alpha[1])+1)>\Gamma_{0}$;
    • when $\alpha>\Gamma_{0}$, by inductive hypothesis we have $\alpha[1]\geq\Gamma_{0}$, and then $\delta[1]=\varphi^{2}_{\delta_{0}[1]}(\varphi_{\delta}(\alpha[1])+1)>\Gamma_{0}$.
  • If $\delta=\Gamma_{\xi}$ with $\xi>0$ then $\delta[1]=\varphi_{\varphi_{\Gamma_{\xi\lceil n\rceil}+1}(0)}(0)>\Gamma_{0}$.∎
@@
  • If $\beta=\varphi_{\xi}(0)$ then $\xi>\gamma$. Moreover $\beta>\xi$ otherwise it would be $\beta=\xi=\varphi_{\xi}(0)$, contradicting the assumption of the Lemma that $\beta$ is not a $\Gamma$ ordinal. Therefore by inductive hypothesis $\xi[1]\geq\gamma$ and so $\beta[1]=\varphi^{2}_{\xi[1]}(0)>\gamma$.
  • If $\beta=\varphi_{\xi}(\alpha)$ with $0<\alpha\leq\gamma$ then it must be $\xi\geq\gamma$ as otherwise $\beta\leq\Gamma_{\gamma_{1}}=\gamma$, a contradiction. Hence $\beta[1]=\varphi^{2}_{\xi[1]}(\varphi_{\xi}(\alpha[1])+1)\geq\varphi^{2}_{\xi[1]}(\varphi_{\gamma}(\alpha[1])+1)>\gamma$.
  • If $\beta=\varphi_{\xi}(\alpha)$ with $\alpha>\gamma$ then, since $\beta>\alpha$, by inductive hypothesis $\alpha[1]\geq\gamma$. Hence $\beta[1]=\varphi^{2}_{\xi[1]}(\varphi_{\xi}(\alpha[1])+1)>\varphi^{2}_{\xi[1]}(\gamma)\geq\gamma$. ∎
@@
Let $\gamma$ and $\delta$ be ordinals such that $\gamma$ is not of the form $\varphi_{\gamma_{1}}(0)$ for any $\gamma_{1}$. If $\gamma>\varphi_{\delta}(0)$ then $\gamma[1]\geq\varphi_{\delta}(0)$ and so $\gamma[n]\geq\varphi_{\delta}(0)$ for each $n>0$.
@@
If $\gamma$ is decomposable or $\delta=0$ the thesis is trivial. We can assume $\gamma=\varphi_{\xi}(\gamma_{0})>\gamma_{0}>0$. Case $\delta=1$ was proved in Lemma 8.2. Case $\xi<\delta$ follows by Lemma 8.3: this is because $\gamma\notin\operatorname{ran}\varphi_{\delta}$ as otherwise $\gamma$ would be a fixed point of $\varphi_{\xi}$, contradicting our assumption that $\gamma>\gamma_{0}$. Finally, if $\xi\geq\delta$ then $\gamma[1]=\varphi^{2}_{\xi[1]}(\varphi_{\xi}(\gamma_{0}[1])+1)\geq\varphi^{2}_{\xi[1]}(\varphi_{\delta}(\gamma_{0}[1])+1)>\varphi_{\delta}(0)$. ∎
@@
first recall that $\gamma[n]=\omega^{\gamma_{0}}+\gamma_{1}[n]$. To prove 1. assume towards a contradiction that $\gamma[n]<\beta[n]<\gamma<\beta$ and consider different subcases depending on the form of $\beta$.
@@
    $\beta=\Gamma_{\xi}$. Let $\delta_{n+1}=\Gamma_{\xi\lceil n\rceil}+1$ and $\delta_{i}=\varphi_{\delta_{i+1}}(0)$ for $i\leq n$. Thus $\Gamma_{\xi}[n]=\varphi_{\delta_{0}}(0)$. Then our assumption $\gamma[n]<\beta[n]<\gamma<\beta$ yields $\gamma_{0}[n]<\delta_{0}<\gamma_{0}<\beta$. We claim that $\gamma_{0}=\varphi_{\gamma_{1}}(0)$ for some $\gamma_{1}$. In fact, if $\gamma_{0}$ were decomposable then, since $\delta_{0}$ is indecomposable, $\gamma_{0}[n]\geq\delta_{0}$, a contradiction. If instead $\gamma_{0}=\varphi_{\gamma_{1}}(\gamma_{2})>\gamma_{2}>0$ then, since $\delta_{0}=\varphi_{\delta_{1}}(0)$, Lemma 8.4 implies $\gamma_{0}[n]\geq\delta_{0}$. This proves the claim. Notice that it must be $\gamma_{0}>\gamma_{1}>0$. The claim and the chain of inequalities $\gamma_{0}[n]<\delta_{0}<\gamma_{0}<\beta$ yields $\gamma_{1}[n]<\delta_{1}<\gamma_{1}<\beta$. Iterating $n+1$ times we obtain for each $i\leq n+1$ an ordinal $\gamma_{i}$ such that $\gamma_{i-1}=\varphi_{\gamma_{i}}(0)$ and $\gamma_{i}[n]<\delta_{i}<\gamma_{i}<\beta$. In particular $\gamma_{n+1}[n]<\delta_{n+1}<\gamma_{n+1}<\beta$. Since $\gamma_{n+1}$ is not a $\Gamma$ ordinal and $\gamma_{n+1}>\delta_{n+1}=\Gamma_{\xi\lceil n\rceil}+1$, Lemma 8.3 implies $\gamma_{n+1}[n]\geq\Gamma_{\xi\lceil n\rceil}$. Since $\Gamma_{\xi\lceil n\rceil}+1>\gamma_{n+1}[n]$ it must be $\gamma_{n+1}[n]=\Gamma_{\xi\lceil n\rceil}$. Lemma 8.5 implies that $\gamma_{n+1}=\Gamma_{\xi\lceil n\rceil}+1$, contradicting $\Gamma_{\xi\lceil n\rceil}+1=\delta_{n+1}<\gamma_{n+1}$.